# PGD-based optimization of 3D bobsleigh track centerlines from 2D centerlines for simulation applications

### v1

@@ 3.2 Tracks Proportionally Scaled from Real 2D Planar Data
Figure 9 presents a comparison between the generated and actual $x$ and yy coordinates of the track centerline. As shown in Figure 9, the $x$ and $y$ coordinates calculated by the proposed algorithm follow the same trend as the actual values. The computed scaling factors $f$ are 1.1102, 1.1196, and 1.0857, corresponding to relative errors of 0.1%, 7.6%, and 2.3%, respectively (with respect to $f$ = 1.1). These results demonstrate that the proposed algorithm can effectively estimate the original $x$ and $y$ coordinates of the track centerline prior to scaling.
Figure 9: Comparison of the track slope variation along the distance for the generated and actual Cesana track centerlines.
To further analyze the computational results, Table 5 presents a comparison between the geometric parameters of the generated and actual track centerlines. As shown in Table 5, the simulated values of track length, height difference, and average slope of the generated track are all close to those of the actual track, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively. Additionally, all slope values remain within the prescribed limits.
Table 5: The comparison of geometric parameters of the generated and actual track centerlines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1291.2 146.9 No 0.1150
Generated track 2 1290.8 143.9 No 0.1126
Generated track 3 1292.2 154.6 No 0.1210
Actual track 1278 149.3 No 0.1160
@@
Figure 10: Comparison of track height variation along the distance for the generated and actual track centerlines.
Figure 11: Comparison of the track slope variation along the distance for the generated and actual track centerlines.
Figures 10 and 11 respectively compare the variation of track height and slope along the distance between the generated and actual track centerlines. Similar to the results in model verification section, the height variation trends of the centerlines generated by the proposed algorithm are consistent with those of the actual track, and all slope values of the generated centerlines remain within the design limits.
In summary, the proposed algorithm can compute the original (pre-scaled) $x$ and $y$ coordinates from the scaled $x–y$ plane coordinates. Similar to the algorithm based on the original unscaled $x–y$ plane coordinates, the one that incorporates the scaling factor is also capable of generating track centerlines with diverse characteristics while maintaining consistency with the overall trends of the actual track centerline. The generated centerlines sufficiently cover the range of the original centerline. Validation using the selected track demonstrates that the generated track’s length, height difference, and average slope are all close to the actual values, with maximum errors of 1.1%, 3.5%, and 4.3%, respectively. Moreover, the slope values do not exceed the specified limits, thereby verifying the effectiveness of the proposed track generation algorithm.
The track generation results are mainly influenced by parameters such as the randomly assigned slope of each segment, the track’s height difference, length, average slope, and maximum slope.
4 Results Analysis: Influence of Different Parameters on Track Generation with Actual 2D Plane Data
In this section, the Whistler and Cesana tracks are selected as case studies to analyze the effects of segment count and weighting factors on track generation under the condition of actual 2D plane data.
4.1 Whistler Track
4.1.1 Influence of Track Segment Count
@@
For the Whistler track, the effect of dividing the track into 15, 26, and 33 segments on the generation results was analyzed. Table 6 presents a comparison of the geometric parameters between the generated and actual Whistler track centerlines. As shown in Table 6, under different segmentation settings, the simulated values of track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively. Additionally, all slope values remain within the specified limits.
Figure 12 shows the comparison of height variation along the distance between the generated and actual track centerlines. It can be observed that the height profiles of the generated centerlines follow the same trend as that of the actual track. For the Whistler track, increasing the number of track segments results in greater local slope variation, but the overall variation remains relatively minor. This is due to the increased number of local slope changes caused by a higher segmentation count. Conversely, when the number of segments is lower, local variations are reduced, while the overall slope variation becomes more pronounced, leading to better agreement with the actual track.
In this case, as the track distance increases, the height also increases. In the 0–600 m segment of the track, the generated centerlines based on 15 and 33 segments show higher consistency with the actual track. In the 600–1200 m segment, the track centerline generated with 15 segments aligns more closely with the actual track. This is because in the 0–600 m segment, the actual centerline exhibits a relatively steep slope, so a larger slope value in the generated centerline leads to a closer match with the actual track. In contrast, in the 600–1200 m segment, the actual centerline has a gentler slope, so a smaller slope value in the generated centerline results in better agreement with the actual track.
@@
Table 6: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=15$ 1291.5 151.4 No 0.1198
Track segment count $s=26$ 1291.7 156.5 No 0.1249
Track segment count $s=33$ 1291.7 156.2 No 0.1222
Actual track 1278 149.3 No 0.1160
Figure 12: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.1.2 Influence of The Weighting Factor
The height difference largely determines the elevation change of each segment along the track. Therefore, this section analyzes the influence of the weighting factor for height difference on the centerline generation results. The number of track segments is set to 26. Table 7 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table, the simulated values of track length, height difference, and average gradient under different height difference weighting factors are all close to the actual values. The maximum errors are 0.7%, 1.7%, and 3.8%, respectively. Furthermore, all slope values remain within the specified limits. As the height difference weighting increases, the height difference of the generated centerline becomes more consistent with the actual value.
Figure 13 shows a comparison between the generated and actual centerline track height variation with distance. As illustrated in Figure 13, the variation trend of the generated centerline height is consistent with that of the actual track. For the segment between 200 m and 600 m, the generated centerlines with weighting factors of 0.7 and 0.9 demonstrate a higher degree of agreement with the actual data. For the segment between 600 m and 1000 m, those with weighting factors of 0.5 and 0.7 are more consistent with the actual centerline.
This is because for similar track lengths, smaller height differences generally correspond to gentler gradients. In the segment between 200 m and 600 m, the generated centerline with a larger gradient shows better agreement with the actual track, whereas in the segment from 600 m to 1000 m, the centerline with a smaller gradient corresponds more closely to the actual values.
Table 7: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1287.6 151.9 No 0.1183
Weight factor $b=0.7$ 1287.6 151.4 No 0.1197
Weight factor $b=0.9$ 1287.5 151.1 No 0.1204
Actual track 1278 149.3 No 0.1160
Figure 13: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.2 Cesana Track
4.2.1 Influence of Track Segment Count
For the Cesana track, the influence of dividing the track into 12, 21, and 29 segments on the centerline generation results is analyzed. Table 8 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table 8, under different numbers of segments, the simulated values of track length, height difference, and average gradient are all close to the actual values, with maximum errors of 0.2%, 0.9%, and 2.4%, respectively. In all cases, the gradient remains within the specified limits.
Figure 14 shows the variation of track height along the distance for both the generated and actual centerlines. As can be seen, the generated centerline height follows a trend consistent with the actual track. For the segments from 0 to 600 m and after 1000 m, the centerline generated with 29 segments exhibits the highest degree of agreement with the actual centerline. In contrast, for the segment from 600 m to 1000 m, the centerline generated with 12 segments shows the highest level of agreement.
This is because increasing the number of segments leads to more local variations in the track gradient, while the overall variation remains limited. Conversely, a smaller number of segments results in fewer local variations but more pronounced overall gradient changes. For this particular track, in the segments from 0 to 600 m and after 1000 m, where the actual centerline has relatively steep gradients, the generated centerline better matches the actual one when the gradient values are higher. In the segment from 600 m to 1000 m, where the actual track gradient is gentler, the generated centerline more closely matches the actual track when the gradient values are smaller.
Table 8: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=12$ 1356.6 123.4 No 0.0932
Track segment count $s=21$ 1356.4 124.0 No 0.0932
Track segment count $s=29$ 1356.6 123.9 No 0.0931
Actual track 1359.2 124.5 No 0.0910
Figure 14: Comparison of track height variation along the distance for the generated and actual track centerlines.
4.2.2 Influence of Weighting Factor
This section also analyzes the influence of the height difference weight factor on the results of centerline generation, with the number of track segments set to 12. Table 9 presents the comparison of geometric parameters between the generated and actual track centerlines. As shown in Table 9, under different weight factors for height difference, the generated track length, height difference, and average slope are all close to those of the actual track, with maximum errors of 8%, 9%, and 5%, respectively, and all slope values remain within the allowable range.
Figure 15 shows the comparison of track height versus distance between the generated and actual track centerlines. As seen in Figure 15, in the section of 0–600 m, the generated track centerline with a weight factor of 0.1 exhibits better agreement with the actual centerline. In the section of 600–1300 m, the generated centerlines with weight factors of 0.7 and 0.9 show better agreement with the actual track.
This is because for similar track distances, a smaller height difference generally results in a smoother profile as indicated by the average slope. In the 0–600 m section, a larger slope in the generated track centerline better matches the actual track, while in the 600–1300 m section, a smaller slope in the generated centerline aligns more closely with the actual track centerline.
Table 9: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.1$ 1357.1 128.9 No 0.0974
Weight factor $b=0.7$ 1356.8 125.0 No 0.0940
Weight factor $b=1.3$ 1356.7 124.7 No 0.0941
Actual track 1359.2 124.5 No 0.0910
Figure 15: Comparison of track height variation along the distance for the generated and actual track centerlines.
To summarize, by dividing the track into different numbers of segments or adjusting the weight of height difference in the optimization equation, multiple centerlines with distinct characteristics can be generated. Within the parameter ranges used in this study, the generated track length, height difference, and average slope under different segmentations are all close to those of the actual tracks, with maximum errors of 1.1%, 4.8%, and 7.7%, respectively. Under different height difference weights, the maximum errors in track length, height difference, and average slope are 0.7%, 3.4%, and 7.0%, respectively. Moreover, all generated slopes remain within the prescribed limits. The agreement between the generated and actual centerlines varies depending on the track and the segmentation or weight factor used. The variety of centerlines generated by the proposed algorithm can be flexibly applied to meet specific track design requirements, thereby serving as a valuable aid in centerline design.
5 Results Analysis: Influence of Different Parameters on Track Generation under Scaled 2D Planar Data
In this section, the Nagano track in Japan and the Pyeongchang track in South Korea are selected as case studies to analyze the influence of segmentation and weight factors on track generation based on scaled 2D planar data.
5.1 Nagano Track
This section selects the Nagano track and the PyeongChang track as case studies to analyze the influence of the number of segments and the weight coefficient on the generated track under conditions involving scaled 2D planar data. Since the starting point of the final ascending segment of the track cannot be determined and the slope design rules for the ascending segment are similar to those for the descending segment bi:14 , the final ascending segment is treated as part of the descending segment. Its height change is included in the total height difference for analysis.
Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Nagano track. The actual geometric parameters of the track are listed in Tables 10 bi:14 .
Table 10: Actual geometric parameters of the Nagano track
@@ 5.1 Nagano Track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1700 118 15 7
5.1.1 Track Generation
Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track in Japan. The calculation results are shown in Table 11. Similar to the results in model verification section, the generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively. Moreover, all slope values remain within the prescribed limits.
Table 11: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1625.8 123.7 No 0.0760
Generated track 2 1652.4 115.6 No 0.0699
Generated track 3 1658 113.9 No 0.0686
Actual track 1700 118 No 0.0700
Figure 16: Comparison of track height variation along the distance for the generated and actual track centerlines.
5.1.2 Influence of Track Segment Count
For the Nagano track, the influence of dividing the track into 10, 15, and 21 segments on the centerline generation results is analyzed. Table 12 presents a comparison of the geometric parameters between the generated and actual track centerlines. As shown Table 12, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 4.4%, 3.6%, and 8.6%, respectively. Furthermore, all slope values remain within the specified limits.
Figure 18 shows the track height variation along the distance for the generated centerlines. Similar to the results of track generation with actual 2D plane data section, Figure 18 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track. A greater number of track segments results in more variations in local slopes while the overall change remains relatively small. This is because increasing the number of segments leads to more pronounced local slope fluctuations. Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
Table 12: Comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=10$ 1653.3 113.8 No 0.0688
Track segment count $s=15$ 1625.8 123.7 No 0.076
Track segment count $s=21$ 1656.4 117.3 No 0.0708
Actual track 1700 118 No 0.07
@@ 5.1.2 Influence of Track Segment Count
Figure 17: Track height variation along the distance for the generated centerlines.
This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 15. Table 13 presents the comparison of the geometric parameters between the generated and actual track centerlines. As shown in Table 13, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 6%, 6%, and 6%, respectively. All slope values remain within the specified limits. As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
Figure 18 shows the track height variation along the distance for the generated centerlines. It can be observed that the height of the generated track centerline changes in accordance with the actual track trend. When the height difference weight is 0.7 or 0.9, the generated track height is close to the actual height. In contrast, when the weight is 0.5, the average slope of the generated track is smaller, resulting in lower track height beyond 600 m compared with those generated with height difference weights of 0.7 and 0.9.
Table 13: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1625.8 123.7 No 0.0760
Weight factor $b=0.7$ 1666.5 115.4 No 0.0692
Weight factor $b=0.9$ 1634.3 117.7 No 0.0713
Actual track 1700 118 No 0.07
@@
Figure 18: Track height variation along the distance for the generated centerlines.
5.2 Pyeongchange Track
Under the algorithm proposed in this study, an analysis was conducted on the proportionally scaled 2D data of the Pyeongchang track. The actual geometric parameters of the track are listed in Table 14 bi:14 .
Table 14: Actual geometric parameters of the Nagano track
Total Length (m) Height Difference (m) Maximum Slope (%) Average Slope (%)
1659 154 25 9.48
5.2.1 Track Generation
Under the proposed algorithm, a track generation analysis is conducted based on the uniformly scaled 2D data of the Nagano track. The calculation results are shown in Table 15. The generated track length, height difference, and average slope are all close to the actual values, with maximum errors of 1.0%, 1.0%, and 6.1%, respectively. Moreover, all slope values remain within the prescribed limits.
Figure 19 shows the track height variation along the distance for the generated centerlines. And the trends and values of different simulation results are similar and close to each other.
Table 15: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Generated track 1 1647.8 152.0 No 0.0921
Generated track 2 1642.3 153.8 No 0.0935
Generated track 3 1641.6 154.0 No 0.0937
Actual track 1659 154 No 0.0981
Figure 19: Track height variation along the distance for the generated centerlines.
5.2.2 Influence of The Track Segment Count
For the Pyeongchang track, the influence of dividing the track into 11, 16, and 22 segments on the centerline generation results is analyzed.
Table 16 presents a comparison of the geometric parameters between the generated and actual track centerlines. As shown in Table 16, the track length, height difference, and average slope of the generated track under different numbers of track segments are all close to those of the actual track, with maximum errors of 1.0%, 4.8%, and 9.8%, respectively. Furthermore, all slope values remain within the specified limits.
Figure 20 shows the track height variation along the distance for the generated centerlines. Similar to the results of track generation in the actual 2D plane data section, Figure 20 demonstrates that the height of the generated centerline varies with distance in accordance with the actual track. A greater number of track segments results in more variations in local slopes while the overall change remains relatively small. This is because increasing the number of segments leads to more pronounced local slope fluctuations. Conversely, with fewer segments, the local variation is less significant, but the overall slope changes more substantially.
Table 16: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Track segment count $s=11$ 1653.6 146.6 No 0.0885
Track segment count $s=16$ 1642.3 153.8 No 0.0935
Track segment count $s=22$ 1667.8 151.6 No 0.0907
Actual track 1659 154 No 0.0981
Figure 20: Track height variation along the distance for the generated centerlines.
5.2.3 Influence of The Weighting Factor
This section also analyzes the effect of the weight factor for height difference on the generated track centerline, with the number of track segments set to 16. Table 17 presents the comparison of the characteristic parameters between the generated and actual track centerlines. As shown in Table 17, the generated track’s length, height difference, and average slope are all close to those of the actual track, with maximum errors of 0.7%, 4.5%, and 10.0%, respectively. All slope values remain within the specified limits. As the weight of the height difference increases, the height difference of the generated centerline becomes closer to that of the actual track.
Figure 21 shows the track height variation along the distance for the generated centerlines. As shown in Figure 21, the height of the generated track centerline varies with track distance in a manner consistent with the trend of the actual track. When the track distance is between 0 and 600 m, the generated track heights under different height difference weights are close. However, after 400 m, due to the smaller average slope of the generated track, the track height generated with a height difference weight of 0.5 is lower than those generated with weights of 0.7 and 0.9.
Table 17: The comparison of geometric parameters of the generated and actual track centrelines
Total Length (m) Height Difference (m) Whether the slope exceeds the limit value Average Slope
Weight factor $b=0.5$ 1662.4 147.0 No 0.0883
Weight factor $b=0.7$ 1647.8 152.0 No 0.0921
Weight factor $b=0.9$ 1650.0 155.3 No 0.0940
Actual track 1659 154.0 No 0.0981
Figure 21: Track height variation along the distance for the generated centerlines.
In summary, the proposed algorithm can effectively generate tracks for two cases with scaled 2D planar data. The generated track length, height difference, and average slope are close to the actual values, with maximum errors of 4.3%, 4.8%, and 8.6%, respectively. The slope values do not exceed the specified limits. On this basis, similar to the results of track generation with actual 2D plane data section, by changing the number of segments or adjusting the weight of height difference in the optimization equation, multiple stylistically different track centerlines can be generated. Within the parameter range used in this study, the maximum errors for track length, height difference, and average slope under different segmentation conditions are 4.4%, 4.8%, and 9.8%, respectively. Under different weight values for height difference, the maximum errors for track length, height difference, and average slope are 4.4%, 4.8%, and 10.0%, respectively.

### v2

@@ 3.2 Tracks Proportionally Scaled from Real 2D Planar Data
Figure 5 presents the comparison between the generated and actual $x$ and $y$ coordinates of the track centerlines. As shown in figure 5, the $x$ and $y$ coordinates calculated by the proposed algorithm follow the same trend as the actual values. The computed scaling factor $f$ is 1.1126, corresponding to a relative error of 0.14% (with respect to $f$ = 1.1111). The result demonstrates that the proposed algorithm can effectively estimate the original $x$ and $y$ coordinates of the track centerline prior to scaling.
@@
Figure 5: The comparison between the generated and actual $x$ and $y$ coordinates of the track centerlines.
@@
To further analyze the computational results, Table 2 presents a comparison between the geometric parameters of the generated and actual track centerlines. As shown in Table 2, Simulation results show that the length, height difference, and average gradient of the generated tracks are close to those of the actual tracks, with maximum deviations of 0.87%, 3.8%, and 3.6%, respectively. Similar to Section 3.1, the generated tracks exhibit smaller average differences between the 2D and 3D curvature compared with their actual counterparts, while the slope values remain within the defined design limits.
@@
Table 2: Comparison of geometric parameters of the generated and actual Whistler track.
@@
Total Length (m) Height Difference (m) Average difference between 2D and 3D curvature Max Slope Average Slope
Generated Track (Whistler) 1289.1 154.9 4.17×10-4 0.1710 0.1218
Actual Track (Whistler) 1278 149.3 9.76×10-4 0.204 0.1176
In summary, the results obtained after applying the scaling factor further demonstrate the robustness and generalizability of the proposed optimization equation. Although minor deviations arise due to the scaling process, the overall geometric characteristics of the generated centerlines remain consistent with those of the actual tracks. The trends of the characteristic parameters—such as height difference, total length, and average gradient—are preserved, and the slopes stay within the defined design limits. These findings indicate that the proposed track centerline generation algorithm maintains accuracy and stability under scaled conditions, thereby confirming its applicability to tracks of varying dimensions. Moreover, the analysis of the scaled cases provides valuable insights for improving the parameter selection and enhancing the adaptability of the algorithm.
4 Influence of Different Parameters on Track Generation
In this section, tracks with real 2D centerline data (Whistler and Cesana tracks) and scaled 2D data (Nagano and Pyeongchang tracks) are selected as case studies to analyze the effects of segment count and weighting factors on track generation under the condition of actual and scaled 2D plane data. And the number of track segments and learning rate are determined separately for each track.
4.1 Influence of the Number of Track Segments
Figure 6 is influence of the number of track segments on track height variation along the track distance. As shown in figure 6, the proposed method reproduces the global geometric characteristics of each track with small deviations, demonstrating stable optimization performance under different segmentation conditions. Simulation results for all four tracks indicate that the maximum deviation ranges for length, height difference, and average gradient are 0.1%-1.1%, 1.3%-13.6% and 1.1%-10.7%, respectively, with detailed values provided in Appendix A (Table 36). All slope values remain within the specified design limits, confirming the robustness of the optimization framework under different weighting conditions.
Increasing the number of segments primarily provides greater local flexibility, allowing more detailed adjustments of the segment-wise gradients. However, the effect of segmentation varies among tracks. For tracks with real 2D centerline data, show greater sensitivity to segmentation, as finer divisions better capture their local slope changes. In contrast, for tracks using scaled 2D data are less affected by segmentation count, showing stable results across different settings.
Overall, these results suggest that the sensitivity to segmentation differs across tracks, and that selecting an appropriate number of track segments can effectively reduce deviations between the generated and actual track data, improving the geometric accuracy of the reconstructed centerlines.
@@ 4.1 Influence of the Number of Track Segments
Figure 6: Influence of the number of track segments on track height variation along the track distance.
4.2 Influence of the Weighting Factor
The weighting factor of the height-difference term determines the degree to which the optimization prioritizes vertical conformity between the generated and actual tracks. In this section, its influence is analyzed across four representative tracks under varying weighting values (0.5, 1.0, and 1.5). Detailed values are provided in Appendix A (Table 710)
Figure 7 is influence of the weighting factor on track height variation along the track distance. As shown in figure 7, Across all cases, the proposed algorithm produces track centerlines whose length, height difference, and average gradient closely match those of the actual tracks, with maximum deviation ranges of 0.2–0.7%, 1.0–16.4%, and 0.8–13.5%, respectively. Similar to the results obtained for the number of track segments, all slope values remain within the specified design limits.
However, the sensitivity to the height-difference weighting varies among tracks. For tracks with real 2D centerline data, increasing the weighting factor enhances the alignment of the generated elevation profile with the actual track, as the optimization places greater emphasis on matching the total height difference. In contrast, for tracks using scaled 2D data, the effect of the weighting factor is less pronounced, since their geometric variation is already constrained by the scaling ratio. Consequently, the influence of the height-difference weighting is track-dependent, and selecting an appropriate weighting factor can effectively reduce deviations between the generated and actual track data, further improving the overall accuracy of the reconstructed 3D centerlines.
@@ 4.2 Influence of the Weighting Factor
Figure 7: Influence of the weighting factor on track height variation along the track distance.